A firm produces candles.the market for candles is highly competitive with candles currently selling for $10.the firm's shortrun total cost function is c=200+0.2q so it's marginal cost is mc=0.4q.what is the firm's profit maximizing quantity?
1
Expert's answer
2017-04-03T11:30:07-0400
For competitive market, the maximum profit corresponds to the condition:
P=MC
From where we can find Q:
P=0.4Q→Q=P/0.4=10/0.4=25